OpenStudy (anonymous):

A rock is thrown vertically with a velocity of 20m/s from the edge of a bridge 42m above a river. How long does the rock stay in the air?

OpenStudy (anonymous):

First Step is to find out how far the rock goes up before dropping, at Ymax V=0 At the point acceleration due to gravity takes over So in order to find Ymax we have to find the time at which it starts to drop Vfinal = 0 So... V=Vo-at 0=Vo-9.8t Solve for t and you have the time it continues to increase in height before dropping That solves for the first half of the duration. Second Step: This is the second half of the duration where it starts to fall and it finally hits the ground meaning Yfinal = 0 but also remember your'e 42m above the river so you had a height advantage to begin with at 42m Now you start at a height of 42m so Yo=42m We plug into the equation: Y-Yo=Vot-(1/2)at^2 Vo is actually zero in this case because it's dropping dropping from Ymax therefore v=0 at max height. 0-42=0-(1/2)at^2 Solving for t should give you the second half of the duration Finally: Add t1 and t2 together you get you're total time.

OpenStudy (anonymous):

For the first part, when it is going up, why is the acceleration of gravity used?

OpenStudy (anonymous):

Disregard the first 4 lines of text, I thought you asking something else and then corrected myself.

OpenStudy (anonymous):

Gravity is used because gravity LOVES to keep you DOWN. Without gravity being used if you jumped up you'd never come back down.

OpenStudy (anonymous):

 Is this how the diagram should be?